There were 840 pumpkins in a pumpkin patch,
but it was difficult for farmer Joe to find the
perfect pumpkin.
• Every third pumpkin was too small.
• Every fourth pumpkin was too green.
• Every fifth pumpkin had a broken stem.
• Every sixth pumpkin had the wrong shape.
How many perfect pumpkins did farmer Joe find
in the pumpkin patch?

Answer by stanbon(75887)

There were 840 pumpkins in a pumpkin patch,
but it was difficult for farmer Joe to find the
perfect pumpkin.
• Every third pumpkin was too small.
• Every fourth pumpkin was too green.
• Every fifth pumpkin had a broken stem.
• Every sixth pumpkin had the wrong shape.
How many perfect pumpkins did farmer Joe find
in the pumpkin patch?
----------------------------
Let the number of pumpkins be "x".
----------------------------
# of perfect pumpkins = (1-1/3-1/5-1/6)x
= (30-10-6-5)x/30
= (9/30)x
= (3/10)x
=================
Ans: (3/10)840 = 252 perfect pumpkins
=========================================
Cheers,
Stan H.